Week 17 - Injector Long MD (Initially from Monday 26/04 08:00 to Thursday 29/04 08:00)
=> DELAYED AND REDUCED BY 1 DAY (due to the SPS vacuum leak) => Tuesday 27/04 08:00 to Thursday 29/04 08:00 (the exact new planning will be defined/updated on Monday 26/04)

- Change of SPS supercycle: LHC_4Inj_FB10860_FT500_2010_V1 => It is the nominal LHC cycle for the 25 ns beam: 451.19 GeV/c (new momentum), 18 BP = 21.6 s

- SPS user = LHC1

- RF MMI target = LHC25NS

- Timing sequence in the CBCM = LHC1-0/1/3(1)

- Work already done on this new cycle (on 13/04/2010):

- Cycle sent to hardware

- The timing table was checked except RF => RF tables timing still to be done

- The hardware compensation on the Main dipole and quadrupole copied from the old cycle

- The settings TT10 copied from LHCFAST1

- The 50 ns beam is already ready in the PS (on 13/04/2010) => The SPS cycle was already prepared (LHC_4Inj_FB7260_FT500_2010_V1), but we will not use it during this first MD

- User timing = LHC2

- Cycle sent to hardware

- The timing table was checked except RF => RF tables timing still to be done

- The hardware compensation on the Main dipole and quadrupole copied from the old cycle

- The settings TT10 copied from LHCFAST1

 

Pre-requisites:

- The interlock threshold on the temperature of the MKE kickers should be ~ 90 C (it should be ~ 70 C when there are extractions) => No need to call Etienne Carlier anymore as it is done in SIS.

- In the past, during (the scrubbing run and) the LONG MDs with LHC beam the ZS were RETRACTED / HV to 0 kV / ION TRAPS ON (the ZS expert was informed). During this MD, ZS experts will participate and the first idea is to keep the ZS in place (i.e. not retracting it).

- The cooling circuit for the septa magnets (circuits 214, 414 and 614) should be switched OFF (by TI from the CCC).

- Block the fast vacuum valves OUT in the injection and extraction areas (LSS1, LSS2, LSS4 and LSS6) => Giovanna Vandoni (TE/VSC, 163269 => ~ 20 min) takes care of this (also after the MD to come back to the initial situation). The list of the 11 valves VVFA and VVFB to be blocked during the MD are:

BA 1 : VVFA 11901

BA 2 : VVFA 21701, VVFB 21877, VVFB 21802

BA 4 : VVFA 41638

BA 6 / TT 60 : VVFA 610213, VVFA 61957, VVFA 61637, VVFA 61460, VVFB 61802, VVFB 61877

- The magnets with the e-cloud monitors are 4 dipoles: MDHW51832, 51836, 51838, 51853, all with 0.12 T (MDHW51832 with 25 A). In addition, there is a C-magnet (MDPH51753 with 28.5 A, in series with 754 to compensate) which contains a sample which can be inserted and removed.

- Comment from Elena Shaposhnikova concerning the 2 coated MBB magnets (MBB.51530 and MBB.51550): One should try and avoid during MDs (all MDs, not only this one) beam losses on these magnets.

 

Monday 26/04 (08:00) until Thursday 29/09 (08:00) - SPS ecloud studies => We will increase the bunch intensity from ~ 2-3 x 1010 p/b to ~ 1.3 x 1011 p/b, the number of bunches from 12 to 72, and the number of batches from 1 to 4 to end up with (hopefully) the nominal parameters shown below

Beam conditions: Nominal LHC beam in SPS

Nb = 1.3 x 1011 p/b (at injection) and ~ 1.15 x 1011 p/b (at top energy)

Transverse emittances: e* < 3 microm

Longitudinal emittance: eL* = 0.35 eVs

Participants for the beam set up: OP, RF, ABP

 

Monday 26/04 (08:00 - 10:00)

Prepare in the PS supercyle 1 LHC25 beam (3 BP) + 1 LHC50 (2 BP), which could be used in the SPS until Tuesday 08:00 => The other 13 BP (out of the 18 BP supercycle) are free for INCA tests in the PS. Between Tuesday 08:00 and Thursday 08:00, 4 LHC25 cycles (= 12 BP) should be reserved in the PS (the rest can be used for INCA tests in the PS)

HW compensation to be checked + HV conditioning for the ZS (B. Balhan)

No access required

 

Monday 26/04 (10:00 - 14:00) => Transverse setting-up

Setting-up of Injection and Capture of the 25 ns LHC beam

Injection (kicker timing adjustment)

Orbit (on all the points of the skeleton) Tune and chromaticity measurement and adjustment

#batches = 1 (12 bunches)

Nb  ~ 2-3 x 1010 p/b

Verify that following measurements work:

 

Monday 26/04 (14:00 - 23:00) => Longitudinal setting-up + transverse dampers

Feedforward and longitudinal feedback setting-up

Check transverse dampers

#batches = 1 (24, then 36, then 48, then 60 and finally 72 bunches) (watch the vacuum!!! and in particular the outgassing of the ZS ion trap => ZS experts will participate to the MD: Should be mainly observations first)

Nb = 1.3 x 1011 p/b => PSB and PS experts should make the necessary adjustments to be able to increase the intensity => Reserve ~ 1/2 h for them to move from low to high intensity

 

Monday 26/04 (23:00) - Tuesday 27/04 (08:00)

Scrubbing run + studies of the 2 coated MBB magnets WITHOUT TOUCHING ANYTHING! The idea is to observe whether the losses reduce with time (i.e. with scrubbing) => The most precise beam loss measurements should be performed

#batches = 1 (72 bunches)

Nb = 1.3 x 1011 p/b

 

Tuesday 27/04 (08:00 - 23:00)

Check that the transverse chromaticities can be reduced

Check that the situation is improved if we inject the same beam but with larger transverse emittances (from the PSB, if the beam is already available) or the 50 ns beam (on the same cycle)

Setting-up for multiple injections (from 2 to 4) => When we go to 2 batches, few h are needed for RF adjustments (T. Bohl and G. Hagmann)

#batches = 1-4 (72 bunches) (the number of batches will be maximized compatibly with the ZS => ZS experts will participate to the MD: Some "solutions" could be tried)

Nb = 1.3 x 1011 p/b

Call BT expert (Etienne Carlier, 160093) to do the fine synchronization of the injection kickers

Ecloud feedback studies (W. Hofle or R. De Maria)

 

Tuesday 27/04 (23:00) to Wednesday 28/04 (08:00)

Scrubbing run + studies of the 2 coated MBB magnets WITHOUT TOUCHING ANYTHING! The idea is to observe whether the losses reduce with time (i.e. with scrubbing) => The most precise beam loss measurements should be performed

#batches = 1-4 (72 bunches) (the number of batches for the night will be maximized compatibly with the ZS)

Nb = 1.3 x 1011 p/b

 

Wednesday 28/04 (08:00 - 23:00)

Check that the transverse chromaticities can be reduced

Check that the situation is improved if we inject the same beam but with larger transverse emittances (from the PSB, if the beam is already available) or the 50 ns beam (on the same cycle)

#batches = 1-4 (72 bunches) (the number of batches will be maximized compatibly with the ZS => ZS experts will participate to the MD: Some "solutions" could be tried)

Nb = 1.3 x 1011 p/b

Ecloud feedback studies (W. Hofle or R. De Maria)

 

Wednesday 28/04 (23:00) to Thursday 29/04 (07:00)

Scrubbing run + studies of the 2 coated MBB magnets WITHOUT TOUCHING ANYTHING! The idea is to observe whether the losses reduce with time (i.e. with scrubbing) => The most precise beam loss measurements should be performed

#batches = 1-4 (72 bunches) (the number of batches for the night will be maximized compatibly with the ZS)

Nb = 1.3 x 1011 p/b

Check in the morning (~ 06:00) that the transverse chromaticities can be reduced

Check in the morning (~ 06:00) that the situation is improved if we inject the same beam but with larger transverse emittances (from the PSB, if the beam is already available) or the 50 ns beam (on the same cycle)

 

Thursday 29/04 (07:00 - 08:00) - Recovery from MD (OP)

Unblock the vacuum valves in the injection and extraction areas (Giovanna Vandoni, TE/VSC, 163269 => ~ 20 min)

Ask TI to restart the cooling circuit for the septa magnets (circuits 214, 414 and 614)

J. Borburgh should verify the conditioning of the septa

L. Ducimetiere and Viliam Senaj should verify the conditioning of the MKDV

- Change of supercycle

- Send beam to physics at 08:00

 


Author: Elias Mtral - AB/ABP
Last update: Wednesday, May 26, 2010 08:17:17 AM
Copyright CERN